# Hybrid Feedback for Three-dimensional Convex Obstacle Avoidance (Extended version)

### v4

@@ IX-D Proof of Theorem 1
Since $\mathbf{0}\in(\mathcal{W}_{r_{a}})^{\circ}$, there exists $\mu_{1}>0$ such that $\mathcal{B}_{\mu_{1}}(\mathbf{0})\cap(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))^{\circ}=\emptyset.$ According to (18), there exists $\mu_{2}>0$ such that $\mathcal{B}_{\mu_{2}}(\mathbf{0})\cap\mathcal{J}_{0}^{\mathcal{W}}=\emptyset$. Additionally, as per (21), there exists $\mu_{3}>0$ such that $\mathcal{B}_{\mu_{3}}(\mathbf{0})\subset\mathcal{J}_{1}^{\mathcal{W}}$. We define the set $\mathcal{S}_{\mu}:=\{\xi\in\mathcal{K}|\mathbf{x}\in\mathcal{B}_{\mu}(\mathbf{0})\},$ where $\mu\in(0,\min\{\mu_{1},\mu_{2},\mu_{3}\}).$ Notice that for all initial conditions $\xi(0,0)\in\mathcal{S}_{\mu}$, the control input, after at most one jump corresponds to the move-to-target mode and it steers the state $\mathbf{x}$ towards the origin according to the control input vector $\mathbf{u}(\xi)=-\kappa_{s}\mathbf{x},\kappa_{s}>0$. Hence, for each $\mu\in(0,\min\{\mu_{1},\mu_{2},\mu_{3}\}),$ the set $\mathcal{S}_{\mu}$ is forward invariant for the hybrid closed-loop system (31).
@@
Consequently, for every $\rho>0$, one can choose $\sigma\in(0,\min\{\mu_{1},\mu_{2},\mu_{3},\rho\})$ such that for all initial conditions $\xi(0,0)$ with $d(\xi(0,0),\mathcal{A})\leq\sigma$, one has $d(\xi(t,j),\mathcal{A})\leq\rho$ for all $(t,j)\in\text{ dom }\xi,$ where $d(\xi,\mathcal{A})^{2}=\underset{(\mathbf{0},\bar{\mathbf{h}},\bar{\mathbf{a}},\bar{m},s)\in\mathcal{A}}{\inf}(\lVert\mathbf{x}\rVert^{2}+\lVert\mathbf{h}-\bar{\mathbf{h}}\rVert^{2}+\lVert\mathbf{a}-\bar{\mathbf{a}}\rVert^{2}+(m-\bar{m})^{2}+(s-\bar{s})^{2})=\lVert\mathbf{x}\rVert^{2}.$ Hence, according to [22, Definition 3.1], the target set $\mathcal{A}$ is stable for the hybrid closed-loop system (31). Next, we proceed to establish the convergence properties of the set $\mathcal{A}$.
@@
Attractivity: We aim to show that for the proposed hybrid closed-loop system (31), the target set $\mathcal{A}$ is globally attractive in the set $\mathcal{K}$ using [22, Defintion 3.1 and Remark 3.5]. In other words, we prove that for all initial conditions $\xi(0,0)\in\overline{\mathcal{F}}\cup\mathcal{J}=\mathcal{K}$, every maximal solution $\xi$ to the hybrid closed-loop system is complete and satisfies
@@
Case 1: $m(0,0)=0$. For the hybrid closed-loop system (31), consider a solution $\xi$ initialized in the move-to-target mode. Let us assume $\xi(t_{0},j_{0})\in\mathcal{F}_{0}$ for some $(t_{0},j_{0})\in\text{ dom }\xi,(t_{0},j_{0})\succeq(0,0).$ If $\xi(t,j)\notin\mathcal{J}_{0}$, for all $(t,j)\succeq(t_{0},j_{0})$, then the control input $\mathbf{u}(\mathbf{x},\mathbf{h},\mathbf{a},0,s)=-\kappa_{s}\mathbf{x}$ will steer the state $\mathbf{x}$ straight towards the origin, where $\kappa_{s}>0$. On the other hand, assume that there exists $(t_{1},j_{1})\succeq(t_{0},j_{0})$ such that $\xi(t_{1},j_{1})\in\mathcal{J}_{0}$. Then, according to (27), the control law switches to the obstacle-avoidance mode. As per (18), it is clear that $\mathbf{x}(t_{1},j_{1})\in\mathcal{J}_{0}^{\mathcal{W}}\cap\mathcal{N}_{\gamma_{s}}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i})),$ for some $i\in\mathbb{I}$ and $d(\mathbf{x}(t_{1},j_{1}),\mathcal{O}_{i})=\beta$ for some $\beta\in[r_{a},r_{a}+\gamma_{s}].$ At this instance, according to (27), the proposed navigation algorithm updates the values of the state variable $\mathbf{h}(t_{1},j_{1}+1)=\mathbf{x}(t_{1},j_{1})$, $\mathbf{a}(t_{1},j_{1}+1)\in\mathbf{A}(\mathbf{x}(t_{1},j_{1}))$, $m(t_{1},j_{1}+1)=1$ and $s(t_{1},j_{1}+1)=s(t_{1},j_{1})+1.$ According to (31), $\mathbf{h}(t_{1},j_{1}+1)=\mathbf{h}(t,j)$, $\mathbf{a}(t_{1},j_{1}+1)=\mathbf{a}(t,j)$ and $m(t_{1},j_{1}+1)=m(t,j)$ for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$, where $I_{j_{1}+1}=\{t|(t,j_{1}+1)\in\text{ dom }\xi\}$. To proceed with the proof, we need the following lemma:
@@
    $\mathbf{x}(t,j)\in\mathcal{P}(\mathbf{h},\mathbf{a})\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$,
@@
See Appendix IX-E. ∎
According to Lemma 4, when a solution $\xi$ to the hybrid closed-loop system (31) evolves in the obstacle-avoidance mode, the state $\xi$ eventually enters in the jump set of the obstacle-avoidance mode $\mathcal{J}_{1}$ (24) and the control law switches to the move-to-target mode.
According to Lemma 4, there exists $(t_{2},j_{1}+1)\succ(t_{1},j_{1}+1)$ with $t_{2}<\infty$ such that $\xi(t_{2},j_{1}+1)\in\mathcal{J}_{1}$. Notice that, according to (21) and (22), one has $\|\mathbf{x}(t_{2},j_{1}+1)\|<\|\mathbf{x}(t_{1},j_{1}+1)\|$. In other words, according to Lemma 4, the proposed navigation algorithm ensures that, at the instance where the control switches from the obstacle-avoidance mode to the move-to-target mode, the origin is closer to the point $\mathbf{x}$ than to the last point where the control switched to the obstacle-avoidance mode. Furthermore, when the control input corresponds to the move-to-target mode, it steers the state $\mathbf{x}$ towards the origin under the influence of control $\mathbf{u}(\xi)=-\kappa_{s}\mathbf{x},\kappa_{s}>0.$ Consequently, given that the workspace $\mathcal{W}$ and the obstacles $\mathcal{O}_{i},i\in\mathbb{I}\setminus\{0\}$, are compact, it can be concluded that the solution $\xi(t,j)$ will contain finite number of jumps and will satisfy (36).
Case 2: $m(0,0)=1.$ For the hybrid closed-loop system (31) consider a solution $\xi$ initialized in the obstacle-avoidance mode. Since $m(0,0)=1$, according to (24), $\xi(0,0)\in\mathcal{J}_{1}$. Therefore, according to (30), the control input switches to the move-to-target mode and $m(0,1)=0.$ One can now use arguments similar to the ones used for case 1 to show that the solution $\xi(t,j)$ will contain finite number of jumps and will satisfy (36).
Hence, the target set $\mathcal{A}$ is globally attractive in the set $\mathcal{K}$ for the proposed hybrid closed-loop system (31). In addition, since the set $\mathcal{A}$ is stable for the hybrid closed-loop system (31), it is globally asymptotically stable in the set $\mathcal{K}$ for the hybrid closed-loop system (31) as per [22, Remark 3.5].
IX-E Proof of Lemma 4
First, we prove that when the control input corresponds to the obstacle-avoidance mode, one has $\mathbf{x}(t,j)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ for all $(t,j)\in(I_{j_{1}+1},j_{1}+1)$. To that end, we make use of Nagumo’s theorem [11, Theorem 4.7] and show that when the control input corresponds to the obstacle-avoidance mode, one has
@@ IX-E Proof of Lemma 4
for all $\mathbf{x}\in\partial\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$. This, combined with the fact that the control input vector $\mathbf{u}(\xi)$ is continuous, when it corresponds to the obstacle-avoidance mode, as stated in Lemma 2, ensures that $\mathbf{x}(t,j)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$.
Note that $\partial\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))=\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})\cup\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$. For all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$, one has $\mathcal{P}_{\geq}(\mathbf{0},\mathbf{x}_{\pi})\subset\mathbf{T}_{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))}(\mathbf{x})$, where $\mathbf{x}_{\pi}=\mathbf{x}-\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})$ with $\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})=\Pi(\mathbf{x},\mathcal{O}_{i})$. Also, since the control input corresponds to the obstacle-avoidance mode, for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$, the control vector (11a) is given by $\mathbf{u}(\xi)=\kappa_{r}\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}$. Since for any $\mathbf{q}\in\mathbb{S}^{2}$, the matrix $\mathbf{P}(\mathbf{q})$ is positive semidefinite, one has $\mathbf{p}^{\top}\mathbf{P}(\mathbf{q})\mathbf{p}\geq 0$ for all $\mathbf{p}\in\mathbb{R}^{3}$. Therefore, for all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$, it is true that $\mathbf{x}_{\pi}\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}\geq 0$. This implies that $\mathbf{u}(\xi)\in\mathcal{P}_{\geq}(\mathbf{0},\mathbf{x}_{\pi})\subset\mathbf{T}_{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))}(\mathbf{x})$ for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$, and condition (37) holds true.
Next, for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, one has $\mathcal{P}_{\leq}(\mathbf{0},\mathbf{x}_{\pi})\subset\mathbf{T}_{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))}(\mathbf{x})$. Also, since the control input corresponds to the obstacle-avoidance mode, for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, the control vector (11a) is given by $\mathbf{u}(\xi)=-\kappa_{r}\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}$. As mentioned earlier, for all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, one has $\mathbf{x}_{\pi}\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}\geq 0$. Therefore, $\mathbf{u}(\xi)\in\mathcal{P}_{\leq}(\mathbf{0},\mathbf{x}_{\pi})\subset\mathbf{T}_{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))}(\mathbf{x})$ for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, and condition (37) holds true. As a result, since $\mathbf{x}(t_{1},j_{1}+1)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, one can conclude that
@@
for all $(t,j)\in(I_{j_{1}+1},j_{1}+1)$.
Next, we show that when the control input corresponds to the obstacle-avoidance mode, $\mathbf{x}(t,j)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$ for all $(t,j)\in(I_{j_{1}+1},j_{1}+1)$. When the control input corresponds to the obstacle-avoidance mode, it is given by $\mathbf{u}(\xi)=\kappa_{r}\mathbf{v}(\mathbf{x},\mathbf{a})$, which, as per (12), can be expressed as a linear combination of the vectors $\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}$ and $\mathbf{R}(\mathbf{a})\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}$. Since $\mathbf{P}(\mathbf{a})$, defined in (13), is an orthogonal projection operator and $\mathbf{0}\in\mathcal{P}(\mathbf{h},\mathbf{a})$, for all $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$, one has $\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}\in\mathcal{P}(\mathbf{h},\mathbf{a})$. Additionally, one can show that $\mathbf{R}(\mathbf{a})\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}\in\mathcal{P}(\mathbf{h},\mathbf{a})$. Therefore, for all $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$, one has $\mathbf{v}(\mathbf{x},\mathbf{a})\in\mathcal{P}(\mathbf{h},\mathbf{a})$. As a result, since $\mathbf{x}(t_{1},j_{1}+1)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$, using (38), one can conclude that
@@
for all $(t,j)\in(I_{j_{1}+1},j_{1}+1)$ and claim 1 in Lemma 4 is satisfied.
Next, we proceed to prove claim 2 in Lemma 4 which states that when $\xi(t_{1},j_{1}+1)\in\mathcal{F}_{1}$ for some $(t_{1},j_{1}+1)\in\text{ dom }\xi$, the control input steers the state $\xi$ to the jump set $\mathcal{J}_{1}$ of the obstacle-avoidance mode in finite time $(t_{2},j_{1}+1)\succ(t_{1},j_{1}+1)$ with $t_{2}<\infty.$

### v5

@@ IX Conclusion
Since $\mathbf{0}\in(\mathcal{W}_{r_{a}})^{\circ}$, there exists $\mu_{1}>0$ such that $\mathcal{B}_{\mu_{1}}(\mathbf{0})\cap(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))^{\circ}=\emptyset.$ According to (24), there exists $\mu_{2}>0$ such that $\mathcal{B}_{\mu_{2}}(\mathbf{0})\cap\mathcal{J}_{0}^{\mathcal{W}}=\emptyset$. Additionally, as per (27) and (29), there exists $\mu_{3}>0$ such that $\mathcal{B}_{\mu_{3}}(\mathbf{0})\subset\mathcal{J}_{1}^{\mathcal{W}}$. We define the set $\mathcal{S}_{\mu}:=\{\xi\in\mathcal{K}|\mathbf{x}\in\mathcal{B}_{\mu}(\mathbf{0})\},$ where $\mu\in(0,\min\{\mu_{1},\mu_{2},\mu_{3}\}).$ Notice that for all initial conditions $\xi(0,0)\in\mathcal{S}_{\mu}$, the control input, after at most one jump corresponds to the move-to-target mode and steers $\mathbf{x}$ towards the origin with the control input vector $\mathbf{u}(\xi)=-\kappa_{s}\mathbf{x},\kappa_{s}>0$. Hence, for each $\mu\in(0,\min\{\mu_{1},\mu_{2},\mu_{3}\}),$ the set $\mathcal{S}_{\mu}$ is forward invariant for the hybrid closed-loop system (38).
@@
Consequently, for every $\rho>0$, one can choose $\sigma\in(0,\min\{\mu_{1},\mu_{2},\mu_{3},\rho\})$ such that for all initial conditions $\xi(0,0)$ with $d(\xi(0,0),\mathcal{A})\leq\sigma$, one has $d(\xi(t,j),\mathcal{A})\leq\rho$ for all $(t,j)\in\text{ dom }\xi,$ where $d(\xi,\mathcal{A})^{2}=\underset{(\mathbf{0},\bar{\mathbf{h}},\bar{\mathbf{a}},\bar{m},\bar{s})\in\mathcal{A}}{\inf}(\lVert\mathbf{x}\rVert^{2}+\lVert\mathbf{h}-\bar{\mathbf{h}}\rVert^{2}+\lVert\mathbf{a}-\bar{\mathbf{a}}\rVert^{2}+(m-\bar{m})^{2}+(s-\bar{s})^{2})=\lVert\mathbf{x}\rVert^{2}.$ Hence, according to [29, Definition 3.1], the target set $\mathcal{A}$ is stable for the hybrid closed-loop system (38). Next, we proceed to establish the convergence properties of the set $\mathcal{A}$.
@@
Attractivity: We aim to show that for the proposed hybrid closed-loop system (38), the target set $\mathcal{A}$ is globally attractive in the set $\mathcal{K}$ using [29, Defintion 3.1 and Remark 3.5]. In other words, we prove that for all initial conditions $\xi(0,0)\in\overline{\mathcal{F}}\cup\mathcal{J}=\mathcal{K}$, every maximal solution $\xi$ to the hybrid closed-loop system is complete and satisfies
@@
Case 1: $m(0,0)=0$. For the hybrid closed-loop system (38), consider a solution $\xi$ initialized in the move-to-target mode. Let us assume $\xi(t_{0},j_{0})\in\mathcal{F}_{0}$ for some $(t_{0},j_{0})\in\text{ dom }\xi,(t_{0},j_{0})\succeq(0,0).$ If $\xi(t,j)\notin\mathcal{J}_{0}$, for all $(t,j)\succeq(t_{0},j_{0})$, then the control input $\mathbf{u}(\mathbf{x},\mathbf{h},\mathbf{a},0,s)=-\kappa_{s}\mathbf{x}$ with $\kappa_{s}>0$ will steer the state $\mathbf{x}$ straight towards the origin. On the other hand, assume that there exists $(t_{1},j_{1})\succeq(t_{0},j_{0})$ such that $\xi(t_{1},j_{1})\in\mathcal{J}_{0}$. Then, according to (34), the control law switches to the obstacle-avoidance mode. As per (24), it is clear that $\mathbf{x}(t_{1},j_{1})\in\mathcal{J}_{0}^{\mathcal{W}}\cap\mathcal{N}_{\gamma_{s}}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i})),$ for some $i\in\mathbb{I}$. At this instance, according to (34), the proposed navigation algorithm updates the values of the state variables $\mathbf{h}(t_{1},j_{1}+1)=\mathbf{x}(t_{1},j_{1})$, $\mathbf{a}(t_{1},j_{1}+1)\in\mathbf{A}(\mathbf{x}(t_{1},j_{1}))$, $m(t_{1},j_{1}+1)=1$ and $s(t_{1},j_{1}+1)=s(t_{1},j_{1})+1.$ According to (38), $\mathbf{h}(t_{1},j_{1}+1)=\mathbf{h}(t,j)$, $\mathbf{a}(t_{1},j_{1}+1)=\mathbf{a}(t,j)$ and $m(t_{1},j_{1}+1)=m(t,j)$ for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$, where $I_{j_{1}+1}=\{t|(t,j_{1}+1)\in\text{ dom }\xi\}$. To proceed with the proof, we need the following lemma:
@@
    $\mathbf{x}(t,j)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$;
@@
See Appendix -F. ∎
According to Lemma 6, there exists $(t_{2},j_{1}+1)\succ(t_{1},j_{1}+1)$ with $t_{2}<\infty$ such that $\xi(t_{2},j_{1}+1)\in\mathcal{J}_{1}$. Notice that, according to (27) and (28), one has $\|\mathbf{x}(t_{2},j_{1}+1)\|<\|\mathbf{x}(t_{1},j_{1}+1)\|$. In other words, according to Lemma 6, the proposed navigation algorithm ensures that, at the instance where the control switches from the obstacle-avoidance mode to the move-to-target mode, the origin is closer to the point $\mathbf{x}$ than to the last point where the control switched to the obstacle-avoidance mode. Furthermore, when the control input corresponds to the move-to-target mode, it steers the state $\mathbf{x}$ towards the origin under the influence of control $\mathbf{u}(\xi)=-\kappa_{s}\mathbf{x},\kappa_{s}>0.$ Consequently, given that the workspace $\mathcal{W}$ and the obstacles $\mathcal{O}_{i},i\in\mathbb{I}\setminus\{0\}$, are compact, it can be concluded that the solution $\xi(t,j)$ will have a finite number of jumps and will satisfy (61).
Case 2: $m(0,0)=1.$ For the hybrid closed-loop system (38), consider a solution $\xi$ initialized in the obstacle-avoidance mode. Since $m(0,0)=1$, according to (31), $\xi(0,0)\in\mathcal{J}_{1}$. Therefore, according to (37), the control input switches to the move-to-target mode and $m(0,1)=0.$ One can now use arguments similar to the ones used for case 1 to show that the solution $\xi(t,j)$ will have a finite number of jumps and will satisfy (61).
Hence, the target set $\mathcal{A}$ is globally attractive in the set $\mathcal{K}$ for the proposed hybrid closed-loop system (38). In addition, since the set $\mathcal{A}$ is stable for the hybrid closed-loop system (38), it is globally asymptotically stable in the set $\mathcal{K}$ for the hybrid closed-loop system (38) as per [29, Remark 3.5].
-F Proof of Lemma 6
First, we prove that when the control input corresponds to the obstacle-avoidance mode, one has $\mathbf{x}(t,j)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$. To that end, we make use of Nagumo’s theorem [13, Theorem 4.7] and show that when the control input corresponds to the obstacle-avoidance mode, one has
@@
for all $\mathbf{x}\in\partial\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$. This, combined with the fact that the control vector trajectory $\mathbf{u}(\xi)$ is continuous, when it corresponds to the obstacle-avoidance mode, as stated in Lemma 3, ensures that $\mathbf{x}(t,j)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$.
Note that $\partial\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))=\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})\cup\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$. For all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$, one has $\mathcal{H}_{\geq}(\mathbf{0},\mathbf{x}_{\pi})=\mathbf{T}_{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))}(\mathbf{x})$, where $\mathbf{x}_{\pi}=\mathbf{x}-\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})$ with $\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})=\Pi(\mathbf{x},\mathcal{O}_{i})$. Also, since the control input corresponds to the obstacle-avoidance mode, for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$, the control vector (12a) is given by $\mathbf{u}(\xi)=\kappa_{r}\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi},\kappa_{r}>0$. Since for any two orthonormal vectors $\mathbf{q},\mathbf{s}\in\mathbb{S}^{n-1}$, the matrix $\mathbf{P}(\mathbf{q},\mathbf{s})$ is positive semidefinite, one has $\mathbf{p}^{\top}\mathbf{P}(\mathbf{q},\mathbf{s})\mathbf{p}\geq 0$ for all $\mathbf{p}\in\mathbb{R}^{n}$. Therefore, for all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$, one has $\mathbf{x}_{\pi}\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}\geq 0$. This implies that $\mathbf{u}(\xi)\in\mathcal{H}_{\geq}(\mathbf{0},\mathbf{x}_{\pi})=\mathbf{T}_{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))}(\mathbf{x})$ for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$, and condition (62) holds true.
Next, for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, one has $\mathcal{H}_{\leq}(\mathbf{0},\mathbf{x}_{\pi})=\mathbf{T}_{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))}(\mathbf{x})$. Also, since the control input corresponds to the obstacle-avoidance mode, for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, the control vector (12a) is given by $\mathbf{u}(\xi)=-\kappa_{r}\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi},\kappa_{r}>0$. As mentioned earlier, for all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, one has $\mathbf{x}_{\pi}\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}\geq 0$. Therefore, $\mathbf{u}(\xi)\in\mathcal{H}_{\leq}(\mathbf{0},\mathbf{x}_{\pi})=\mathbf{T}_{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))}(\mathbf{x})$ for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, and condition (62) holds true. As a result, since $\mathbf{x}(t_{1},j_{1}+1)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, one can conclude that
@@
for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$.
Next, we show that when the control input corresponds to the obstacle-avoidance mode, $\mathbf{x}(t,j)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$ for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$. When the control input corresponds to the obstacle-avoidance mode, it is given by $\mathbf{u}(\xi)=\kappa_{r}\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})$, which, as per (13), can be expressed as a linear combination of the vectors $\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}$ and $\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}$. Note that, according to (34), one has $\mathbf{h}^{\top}\mathbf{a}=0$. Therefore, for all $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$, as per (14), one has $\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}\in\mathcal{P}(\mathbf{h},\mathbf{a})$. Additionally, using (15), one can show that $\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}\in\mathcal{P}(\mathbf{h},\mathbf{a})$. Therefore, for all $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$, one has $\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})\in\mathcal{P}(\mathbf{h},\mathbf{a})$. As a result, since $\mathbf{x}(t_{1},j_{1}+1)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$, using (63), one can conclude that
@@
for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$ and claim 1 in Lemma 6 is satisfied.
Next, we proceed to prove claim 2 in Lemma 6 which states that when $\xi(t_{1},j_{1}+1)\in\mathcal{F}_{1}$ for some $(t_{1},j_{1}+1)\in\text{ dom }\xi$, the control input steers the state $\xi$ to the jump set of the obstacle-avoidance mode $\mathcal{J}_{1}$ in finite time $(t_{2},j_{1}+1)\succ(t_{1},j_{1}+1)$ with $t_{2}<\infty.$